OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 15 and a13= −57? −171 −165 −159 −151

OpenStudy (anonymous):

@jdoe0001 i know you can figure this out xD

OpenStudy (anonymous):

Do you know this ? @Peaches15

OpenStudy (anonymous):

Answer is -171 Okay for a general term, T(n) the formula for an a.p. is T(n) = a + (n-1)d where a = first term, n =nth term, d = common difference Since a1 = 15 I assume as in the first term T(1) = 15 thus sub it into the general formula, T(1) = a + (1-1)d = 15 a = 15 since 0 x d gives 0 Okay now that you know a = 15, sub it into the 13th term T(13) = -57 thus -57 = a + (13 - 1)d since you found out that a = 15 earlier you get -57 = 15 + 12d ----> d = -6 so substitute that into the general formula to get the 32nd term T(32) = a + (n-1)d since a = 15 (found from the first part), n = 32 (32nd term) and d = -6 (calculated above) T(32) = 15 + 31(-6) = -171

OpenStudy (anonymous):

@LexiLuvv2431 do I get a medal

OpenStudy (anonymous):

yess thank youu